# Quantformer: from attention to profit with a quantitative transformer trading strategy

### v2

@@ 2.3 Transformer in quantitative trading
Transformer [10], which was introduced in 2017, has received considerable attention in the NLP field. According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models. The transformer-based models such as GPT [34] and BERT [35] are the basis of the specific field of application. In the application field, transformer-based models such as ChatGPT [36], FinBERT [37, 38] and BloomberGPT [39] show advantages and ability in solving different types of problems in different fields. [40] utilized transformer to to predict the stock market index (including CSI 300, S&P 500, Hang Seng Index, and Nikkei 225) and concluded that the model could better catch the rules of stock market dynamics.
Similar to other neural sequence transduction models, transformer also has an encoder-decoder structure. The model takes as input a sequence of context which could be natural language sentences and generates output with texts depending on the input, similar to a translation according to the input or a predicted sentence. Transformer using stacks which are called self-attention and point-wise, which connect layers in the architecture for encoder and decoder. For these encoder and decoder layers, the self-attention method plays an important role in matrix calculation. The linear and softmax functions are used in the dot-product attention module. In the vanilla transformer model, multi-head attention layers are used and the result from each layer will be outputted after going through an additional normalization layer. The above-mentioned steps cover most of the progress in the vanilla transformer.
In the quantitative trading domain, some previous researchers tried to test the potential of transformer. [41] firstly tried to exploit transformer on trading sequences to classify stock price movements. Besides using a traditional transformer, other innovative methods of transformer are worth mentioning. [42] devoted to optimizing the efficiency of time complexity and memory usage of transformer on extremely long time series by informer. [43] takes the advantage of CNNs and transformers to model short-term and long-term dependencies in financial time series. They showed the advantage of the proposed approach on intraday stock price prediction of S&P 500 constituents. [44] proposed a new ConvLSTM model by combining convolutional transformer and LSTM. The model was trained by stocks from S&P 500 from 2004-2021 and performs better than other models such as VaR and ARIMA.
Similar to previous work, [1] innovated the structure of a transformer, in the form of Adaptive Long-Short Pattern Transformer (ALSP-TF). Their model is structurally innovated for hierarchical representation and interaction of stock price series at different context scales. With the help of a learnable function, they make the self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching. In the end, they obtained more than 10% of annual return on average.
However, transformer has some disadvantages. [45] mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights. As fluctuations of stocks are conditioned on composite signals over manifold periods, lacking pattern-wise interaction hinders the adequate discrimination of stock tendency and is susceptible to noise points. On the other hand, [46] claimed that the basic query-key matching paradigm is position agnostic. Although position embedding is inserted into the sequential inputs, it may not be optimal because of the inability to reveal precise distances.
@@ 3 Methodology
This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, quantformer construction and prediction.
@@ 3.1 Framework overview
(ii) Embedding the training dataset and training them by a quantformer with word-embedding layer replaced by linear layer and with no Mask(opt.) layer;
@@ 3.2 Problem formulation
Inputs
Given the candidate stock set $S^{t}$ for $N$ stocks on the trading timestamp $t$:
$S^{t}=\{s_{1}^{t},s_{2}^{t}\ldots s_{N}^{t}\}$
For a stock $s_{k}^{t}$ in the stock set $S^{t}$, where $k\in\{1,\ldots,N\}$, consider the two-dimensional historical 1-time step sequence (such as one month):
${\chi}^{t}_{k}=[x_{k}^{t-20},x_{k}^{t-19}\ldots x_{k}^{t}]\in\mathbb{R}^{20\times 2}$
Each $x_{k}^{t}$ contains two features. The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ are the close price of the $i^{\mathrm{th}}$ and ${(i-1)}^{\mathrm{th}}$ day, respectively.
$r^{t}=\sum r_{i}=\sum\frac{p_{i}-p_{i-1}}{p_{i-1}},\quad\sigma^{t}=\sum\sigma_{i}$ (1)
The other feature is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
$\chi^{t}_{k}=\left[\begin{matrix}r^{t-20}&r^{t-19}&\ldots&r^{t}\\ \sigma^{t-20}&\sigma^{t-19}&\ldots&\sigma^{t}\\ \end{matrix}\right]\in\mathbb{R}^{20\times 2}$
Outputs
For a stock set $S^{t}$ on the trading timestamp $t$, each stock $s^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is the same as for $r^{t}$ in the sequence of inputs shown in equation (1). Then the value in the list of next-time stamp’s profit for $N$ stocks in the timestamps $t+1$ can be represented as:
$P^{t+1}=\{r_{1}^{t+1},r_{2}^{t+1}\ldots r_{N}^{t+1}\}$
This list is then sorted and partitioned into $q$ equal parts, where it is set to 3 or more. For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively. And the dimension of the outputs $d_{f}$ is the number of selected quantiles in training. The value of $d_{f}$ should be larger than 3 and not larger than $q$.
Figure 4: Training outputs remarked between different $d_{f}$
Both the dimensions of training and predicted outputs are the same, which is the $d_{f}$ set at the beginning. For example, set $q=5$; for the stocks $s_{k}$ whose profit values ranked in the bottom $20\%$, middle $20\%$, and top $20\%$ respectively, and the corresponding output of $d_{f}=3$ is recorded, respectively, as
$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
The dimension of the output also can be 5, and in this case, the outputs of $d_{f}=5$ are:
$y^{t}_{k}=[1,0,0,0,0]^{T},\quad y^{t}_{k}=[0,0,1,0,0]^{T},\quad y^{t}_{k}=[0,0,0,0,1]^{T}$
In this example, when $d_{f}=3$, the output dimension will be 3 while the dimension will be 5 as $d_{f}$ is set as 5.
3.3 Embedding initialization
@@
The next step is to normalize the input data for each time step by zero-mean, unit-variance normalization (Z-score), the normalization equation is shown in equation (2).
$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{\mathrm{th}}$ day of stock $k$ at time $t$. $\mathbb{E}[x^{t}_{i}]$ and $\mathrm{std}[x^{t}_{i}]$ are the mean and standard deviation of the variables $x^{t}_{i}$, respectively.
$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-\mathbb{E}[x^{t}_{i}]}{\mathrm{std}[x^{t}_{i}]}$ (2)
In this way, the input sequences are normalized with zero mean and unit variance, which aims to reduce the influence from outlying time points and allow different features to be comparable with each other [47].
@@ 3.4 Quantformer encoder
Then a quantformer encoder structure will be built in this subsection. The representation hierarchy consists of $L$ blocks of multi-head self-attention layers. Taking initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, the canonical self-attention in [10] can perform information exchange between every time points for each stock $s_{i}$.
To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding. On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value representing the probability of price increase or decrease. Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations.
More specifically, each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$ is transformed into query, key, and value metrics, respectively:
$\textbf{Q}_{i,h}=\textbf{P}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{P}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{P}_{i}\textbf{W}^{V}_{h}$ (3)
where $h=1,\ldots,H$ is the head index, as well as trainable weights $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$. Then, the final layer is represented by the concatenation of all attention heads.
$\textbf{F}_{i}=\mathrm{Multihead}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\big{(}||^{h=H}_{h=1}\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})\big{)}\textbf{W}^{O}$ (4)
where ||||| | represents the concatenation operator and $d_{f}$ is the dimension of the projected feature space. After the attention computation is completed, the multi-head attention output is divided into different ”heads”. These outputs from the heads need to be recombined back into the original input dimension. $\textbf{W}^{O}$ $\in\mathbb{R}^{Hd_{f}\times 2}$ in equation 5 plays the role of combining these separate heads back into the original dimension. The attention head is shown below:
$\mathrm{Attention}(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=\mathrm{softmax}(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h}$ (5)
In the end, all input sequences are represented in the form of:
$\textbf{F}=[\textbf{F}_{1};\textbf{F}_{2};\ldots\textbf{F}_{N}]\in\mathbb{R}^{N\times 20\times Hd_{f}}$
which are fed into feed-forward layers. From Section 3.2, the embedding output is denoted as:
$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$
@@ 3.5 Output process
Eventually, the output is represented as:
$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times d_{f}}$ (6)
The output layer uses the softmax activation function, and for each $z^{t}_{i}$ can be written as:
$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,d_{f}}]\in\mathbb{R}^{1\times d_{f}}$
($m=1,2,\ldots,n$ is the index number between $1$ and $n$). The output probability can be represented as:
$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,d_{f}}]^{T}=\left[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{d_{f}}}}\right]$ (7)
In equation (7), the $y^{t}_{o,i}$ take values between 00 and $1$ and sum to $1$, which means that they can be interpreted as the probability of the stock’s performance in the slice $m$ in the next timestamp

### v3

@@ 2.4 Transformer in quantitative trading
In the quantitative trading domain, a few research articles have investigated the potential of transformers. Wang et al. (2022a) utilized transformer to predict stock market indices (including CSI 300, S&P 500, Hang Seng Index, and Nikkei 225) and concluded that the model can better catch the rules of stock market dynamics. Ding et al. (2020) exploited transformer on trading sequences to classify stock price movements. Besides using a traditional transformer, other innovative methods of transformer are worth mentioning. Zhou et al. (2021) optimized the efficiency of time complexity and memory usage of transformer on extremely long time series by informer. Zeng et al. (2023b) combined the advantages of CNNs and transformers to model short-term and long-term dependencies in financial time series. They illustrate the merits of this approach on intraday stock price prediction of S&P 500 constituents. ConvLSTM (Kim et al., 2024) has been trained on stocks from the S&P 500 between 2004 and 2021, and performs better than other models such as VAR and ARIMA.
Similar to previous work, Wang et al. (2022b) introduced Adaptive Long-Short Pattern Transformer (ALSP-TF). Their model is structurally designed for hierarchical representation and interaction of stock price series at different context scales. With the help of a learnable function, they make self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching. In the end, they obtained more than 10% of annual return on average.
However, the transformer model can also have some disadvantages in practice. Xu et al. (2021) mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights. As fluctuations of stocks are conditioned on composite signals over manifold periods, lacking pattern-wise interaction hinders the adequate discrimination of stock tendency and is susceptible to noise points. On the other hand, Wu et al. (2021a) claimed that the basic query-key matching paradigm is position agnostic. Although position embedding is inserted into the sequential inputs, it may not be optimal because of the inability to reveal precise distances.
@@ 3 Methodology
Several researchers have tried to improve transformers for time series and quantitative financial applications and proposed improved models (see time series papers in Table 1), but there are still some aspects that have not been considered. For data in time series type, such as stock prices or daily returns, it is reasonable to assume that each data in the series is assigned a “position” automatically, which is different from sentences. Based on this observation, the positional encoding module on the vanilla model may not be a necessary part. Furthermore, market data are related to the market trend. For example, the return of the stock may include the market sentiment and the turnover rate reflects investor sentiment and trading enthusiasm. Therefore, integrating such market-related features as input can enhance the model’s ability to capture underlying market dynamics and improve the accuracy of future trend predictions. To implement these ideas, we propose quantformer, an enhanced transformer-based model for quantitative financial trading, and assess its performance on real-world datasets.
This section introduces the framework of the work and the steps of the establishment of the model, including data processing, quantformer construction and prediction.
@@ 3.1 Framework overview
(ii) Unlike the transformer, embedding the training dataset and training them by a quantformer with word-embedding layer replaced by linear layer and with no Mask(opt.) layer;
@@ 3.2 Problem formulation
Training input
Consider a candidate stock set $S^{t}$ containing $N$ stocks on the trading timestamp $t$:
$S^{t}=\{s^{t}_{n}\}^{N}_{n=1}$
For a stock $s_{n}^{t}$ in the stock set $S^{t}$, where $n\in\{1,\ldots,N\}$, consider the two-dimensional historical 1-time step (such as one month) input feature matrix $\mathcal{X}^{t}_{n}\in\mathbb{R}^{20\times 2}$, consisting of 20 continuous timestamps:
${\mathcal{X}}^{t}_{n}=\{x^{t-m}_{n}|m=19,18,\dots,0\}$ (1)
Each row vector $x_{n}^{t-m}$ for $m=0,\ldots,19$ contains two features. The first one is the accumulated daily profit rate $r^{t-m}_{n}$ during the time step, where $p^{t-m}_{n}$ is the close price of the last day at the timestamp tmt-m and $p^{t-m-1}_{n}$ is the close price of the last day at the timestamp tm1t-m-1.
$r^{t-m}_{n}=\frac{p^{t-m}_{n}-p^{t-m-1}_{n}}{p^{t-m-1}_{n}},\quad v^{t-m}_{n}=\sum v^{t-m,i}_{n}$ (2)
The second feature is the accumulated daily turnover rate $v^{t-m}_{n}$ during the time step tmt-m for stock $n$. $v^{t-m,i}_{n}$ is the daily turnover rate and $i$ is the day index in the time period. To sum up, each $\mathcal{X}^{t}_{n}$ can be represented as:
$\mathcal{X}^{t}_{n}=\left[\begin{matrix}x^{t-19}_{n}&x^{t-18}_{n}&\ldots&x^{t}_{n}\\ \end{matrix}\right]^{T}\in\mathbb{R}^{20\times 2},x^{t-m}_{n}=[r^{t-m}_{n},v^{t-m}_{n}]$ (3)
@@
The next step is to normalize the input data for each time step by zero-mean, unit-variance normalization (Z-score). The normalization equation is shown in equation (4). For the row vector $x^{t}_{n}$ (equation (1)), $\mathbb{E}[x^{t}]$ and $\mathrm{std}[x^{t}]$ represent the mean and standard deviation of all the row vectors $x^{t}$ at time $t$, respectively, and the resulting values $\tilde{x}^{t}_{n}$ represent the normalized values of the row data $x^{t}_{n}$.
$\tilde{x}^{t}_{n}=\frac{x^{t}_{n}-\mathbb{E}[x^{t}]}{\mathrm{std}[x^{t}]}$ (4)
So, the normalized two-dimensional historical 1-time step input feature matrix can be represented in this way:
${\tilde{\mathcal{X}}}^{t}_{n}=\{\tilde{x}^{t-m}_{n}|m=19,18,\dots,0\}$ (5)
In this way, the input sequences are normalized with zero mean and unit variance, which aims to reduce the influence from outlying time points and allow different features to be comparable with each other (Klambauer et al., 2017).
Target labels
For a stock set $S^{t}$ on the trading timestamp $t$, each stock $s^{t}_{n}$ has a profit rate $r_{n}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is the same as for $r^{t}_{n}$ in the sequence of inputs shown in equation (2). The set $\mathcal{G}^{t+1}$ contains the next-time stamp’s profit for $N$ stocks in the timestamps $t+1$. It is defined as $\mathcal{G}^{t+1}=\{r^{t+1}\}^{N}_{n=1}$. Each $r^{t+1}_{n}$ in the list $\mathcal{G}^{t+1}$ is then ranked by $\Psi(r^{t+1}_{n})$, which is the empirical quantile CDF for $r^{t}_{n}$:
$\Psi(r^{t+1}_{n})=\frac{1}{n}\sum^{n}_{i=1}\mathbbm{1}_{\{r^{t+1}_{i}\leq r^{t+1}_{n}\}}$ (6)
For the target label for each feature $s^{t}_{n}$ in $\mathcal{G}^{t+1}$, we have a one-hot vector $y^{t}_{n}\in\mathbb{R}^{\varrho}$ based on $\Psi(r^{t+1}_{n})$. $y^{t}_{n}$ is partitioned into $\varrho$ equal bins, where $\varrho$ is set to 3 or more. Each bin corresponds to a $\varphi\times 100\%$ of the stocks in the set, where $\varphi\varrho\leq 1$. And the boundary term $\xi$ is used to ensure non-overlapping intervals. Then $y^{t}_{n}$ is represented as:
$y^{t}_{n}=[\mathbbm{1}_{\{(i-1)(\varrho+\xi)\leq\Psi(r^{t+1}_{n})<i\varrho\ +(i-1)\xi\}}|i=1,\dots,\varrho]\in\mathbb{R}^{\varrho},\quad\xi=\frac{1-\varphi\varrho}{\varrho-1}$
For the empirical quantile CDF $\Psi(r^{t+1}_{n})$ of the stock in the range $i\varphi,(i+1)\xi)$ for $i=1,\ldots,\varrho$, $y^{t}_{n}$ is represented as $\mathbf{0}_{\varrho}$. For example, suppose $\varrho=3$, $\varphi=0.2$; for the stocks $s^{t}_{n}$ whose profit values $r^{t}_{n}$ ranked in the top, middle, bottom $\varphi\times 100\%=20\%$ respectively, as ($\xi=0.2$ in this case):
$y_{n}^{t}=\begin{cases}[1,0,0]^{T}&\text{if }\Psi(r_{n}^{t+1})\in[0,0.2)\\ [0,1,0]^{T}&\text{if }\Psi(r_{n}^{t+1})\in[0.4,0.6)\\ [0,0,1]^{T}&\text{if }\Psi(r_{n}^{t+1})\in[0.8,1.0)\\ [0,0,0]^{T}&\text{otherwise}\end{cases}$
The stocks in the intermediate ranges ($\Psi(r^{t+1}_{n})\in[0.2,0.4),[0.6,0.8)$) are marked as $0,0,0]^{T}$. Alternatively, when $\varrho=5$ and $\varphi=0.2$, the stocks are divided into five equally sized quantiles, with no null labels ($0,0,0,0,0]^{T}$), where each part contains $20\%$ of the stock in $S^{t}$:
$y^{t}_{n}\in\{[1,0,0,0,0]^{T},[0,1,0,0,0]^{T},\ldots,[0,0,0,0,1]^{T}\}$
@@ 3.3 Quantformer encoder
Then, a quantformer encoder structure is designed as described in this subsection. The representation hierarchy consists of $L$ blocks of multi-head self-attention layers. Taking initialized stock embedding sequences $\mathcal{X}=\{\tilde{\mathcal{X}}^{(t)}_{n}\}^{N}_{n=1}\in\mathbb{R}^{N\times 20\times 2}$ as inputs, the canonical self-attention layers introduced in Vaswani et al. (2017) can perform information exchange between every time points for each stock $s^{t}_{n}$.
To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding. The linear embedding is shown in equation 7, where $\mathbf{W}_{E}\in\mathbb{R}^{2\times d}$ is the trainable weight matrix and $\theta_{E}\in\mathbb{R}^{d}$ is the bias and $d$ is the dimension of the hidden feature space. The resulting sequence $\mathcal{X}^{\prime}_{i}$ is
$\mathcal{X}^{\prime}_{i}=\mathcal{X}_{i}\mathbf{W}_{E}+\theta_{E}$ (7)
In stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value representing the probability of price increase or decrease. Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations. For the attention head $h=1,\ldots,H$, the query, key, and value metrics for each $\mathcal{X}_{i}\in\mathbb{R}^{20\times 2}$ are computed using separate learned linear projections:
$\mathbf{Q}_{i,h}=\mathcal{X}^{\prime}_{i}\mathbf{W}^{Q}_{h},\quad\mathbf{K}_{i,h}=\mathcal{X}^{\prime}_{i}\mathbf{W}^{K}_{h},\quad\mathbf{V}_{i,h}=\mathcal{X}^{\prime}_{i}\mathbf{W}^{V}_{h}$ (8)
where the trainable weights $\mathbf{W}^{Q}_{h}$, $\mathbf{W}^{Q}_{h}$, $\mathbf{W}^{Q}_{h}$ $\in\mathbb{R}^{d\times\varrho}$ for the $h\mathrm{th}$ head. Then the attention parameter for each head is computed using scaled dot-produced attention:
$\mathrm{Attention}(\mathbf{Q}_{i,h},\mathbf{K}_{i,h},\mathbf{V}_{i,h})=\mathrm{softmax}(\mathbf{Q}_{i,h}\mathbf{K}_{i,h}^{T}/\sqrt{d})\mathbf{V}_{i,h}$ (9)
The outputs from all heads are concatenated and sent back to the original dimension using $\mathbf{W}^{O}\in\mathbb{R}^{\varrho H\times d}$, and the final multi-head attention output $\mathbf{F}_{i}$:
$\mathbf{F}_{i}=\mathrm{Multihead}(\mathbf{Q}_{i,h},\mathbf{K}_{i,h},\mathbf{V}_{i,h})=\big{(}||^{h=H}_{h=1}\mathrm{Attention}(\mathbf{Q}_{i,h},\mathbf{K}_{i,h},\mathbf{V}_{i,h})\big{)}\mathbf{W}^{O}$ (10)
where $||$ represents the concatenation operator. After applying attention to all inputs, all the output from attention modules are represented in the form of:
$\mathbf{F}=[\mathbf{F}_{1};\mathbf{F}_{2};\ldots\mathbf{F}_{N}]\in\mathbb{R}^{N\times 20\times\varrho H}$
which are fed into feed-forward layers.
@@ 3.4 Output process
After processing through the multi-head self-attention and feed-forward layers, with the encoder output $F_{i}$, the output layer can be represented as:
$\mathbf{Z}^{t}=\mathrm{Softmax}(\mathbf{F}_{i}\mathbf{W}_{Z}+\theta_{Z})=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{\varrho}$ (11)
where $\mathbf{W}_{Z}\in\mathbb{R}^{\varrho\times d}$ and $\theta_{Z}\in\mathbb{R}^{d}$. With the softmax function, the predicted probability of the output can be represented as:
$\hat{Y}^{t}_{n}=\{\hat{y}^{t}_{n,i}\}^{\varrho}_{i=1}=\left\{\frac{\exp(z^{t}_{i})}{\sum^{\varrho}_{i=1}\exp(z^{t}_{i})}\right\}^{\varrho}_{i=1}\in\mathbb{R}^{\varrho}$ (12)
In equation (12), the $\hat{Y}^{t}_{n}$ take values between 0 and $1$ and sum to $1$, which means that they can be interpreted as the probability distribution of the stock’s performance.